OpenStudy (cocoacup):

Simplify (x + 4)(x – 3) show your work

OpenStudy (bluebeta):

x*x= x*-3= 4*x= 4*-3= figure that out first

OpenStudy (cocoacup):

-12

OpenStudy (will.h):

(x + 4)(x – 3) x * x = x^2 x * -3 = -3x 4 * x = 4x 4 * -3 = -12 so put all those at once x^2 -3x + 4x - 12 simplify alike terms x^2 + x - 12 questions?

OpenStudy (cocoacup):

Thank you ^^

OpenStudy (will.h):

you welcome let me know if you have additional questions :)